OpenStudy (anonymous):

how do i solve this equation [(-5xy)^2]^5

OpenStudy (lgbasallote):

hint \[\huge [(ab)^m]^n \implies (ab)^{mn}\] does that help?

OpenStudy (anonymous):

yea

OpenStudy (lgbasallote):

wonderful

OpenStudy (anonymous):

two things what do i do with the exponents do i multiply and apply it too the parenthesis and what do i do the variables

OpenStudy (lgbasallote):

first...multiply the outside variables

OpenStudy (lgbasallote):

after you've done that.. "distribute" the exponents to EACH of the variables/numbers in the parenthesis

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i ended up with -125x25y25

OpenStudy (lgbasallote):

what did you do?

OpenStudy (anonymous):

i thought i had to multiply the exponents and distribute it to whats on the inside of the parenthesis

OpenStudy (lgbasallote):

you multiply the exponents right? \[\huge [(-5xy)^2]^5 \implies (-5xy)^{2\times 5} \implies (-5xy)^{10}\] right?

OpenStudy (anonymous):

opps my bad i put 25 instead of ten

OpenStudy (lgbasallote):

by "distribute" i meant like this \[\huge (ab)^m \implies a^m b^m\]

OpenStudy (anonymous):

yes i got that did that but do i also multiply the coefficient by th exponent

OpenStudy (amistre64):

(5n)^3 = 5n*5n*5n = 5*5*5 * n*n*n = 5^3 n^3